40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Fri Jan 22 11:33:20 2016, gautam, Update, Green Locking, AUX-X AM/PM investigations AUX_X_AM.pdfAUX_X_RIN.pdfNE_Mephisto.png
    Reply  Fri Jan 22 13:33:37 2016, ericq, Update, Green Locking, AUX-X AM/PM investigations 
       Reply  Fri Jan 22 17:22:06 2016, gautam, Update, Green Locking, AUX-X AM/PM investigations AM_response.pdfNE_investigations.pdf
          Reply  Fri Jan 22 18:46:03 2016, rana, Update, Green Locking, AUX-X AM/PM investigations 
Message ID: 11945     Entry time: Fri Jan 22 13:33:37 2016     In reply to: 11944     Reply to this: 11946
Author: ericq 
Type: Update 
Category: Green Locking 
Subject: AUX-X AM/PM investigations 
Quote:

Attachment #1 shows the measured AM response. It differs qualitatively in shape from the earlier measurements reported in this elog and on the wiki below the 100kHz region. 

It looks like some of the features may have shifted in frequency. The previous measurement results can be found in /users/OLD/mott/PZT/2NPRO, can you plot the two AM measurements together?

ELOG V3.1.3-